please help, show work.

5.) Jennifer has been given money for her birthday. She puts her money in a savings account offers an annual interest rate of 4.50%, compounded monthly. What is the effective annual rate (EAR) on the saving account?

6.) An individual invests $10,000 today in an investment that is expected to be worth $20,000 in 6 years. What annualized rate of return is the investor expecting to receive on the investment?

7.) A college graduate just bought a new house. The home was purchased with a 30-year loan of $350,000 with monthly payments at an interest rate of 4%. What is the amount of the monthly loan payment?

8.) An investor is evaluating an investment that pays $50,000 for 3 consecutive years and then $25,000 in year 4. The first $50,000 payment comes one year from today. What is the present value of the investment’s cash flows at an interest rate of 10%?
Answer #1

5: EAR =(1+APR/m)^m = (1+4.5%/12)^12-1

=4.59%

6: Using financial calculator

Input: PV = -10000, N = 6, FV=20000

Solve for I/Y as 12.25

Hence rate = 12.25%

7: Using financial calculator

Input: PV = -350000, N=30*12=360 , I/Y = 4/12

Solve for PMT as 1,670.95

amount of the monthly loan payment= 1670.95

8: PV= C0+ CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n

= 50000/1.1^1+50000/1.1^2+50000/1.1^3+25000/1.1^4

= 141417.94